In an RLC series circuit, the source has a potential difference of 160 V (rms). The impedance is 175 Ω and the resistance is 40 Ω . Find:

a) the average power delivered by the source; P = W

b) the power factor;
Answer #1

1) power delivered by the source :

P = (Vrms) ^2*Pf/Z

P = 160*160*0.228/175

P = 33.35 Watt

2) power factor :

Pf = R/Z

Pf = 40/175

Pf = 0.228